OpenStudy (anonymous):

A fruit company delivers its fruit in two types of boxes: large and small. A delivery of 2 large boxes and 3 small boxes has a total weight of 56 kilograms. A delivery of 4 large boxes and 8 small boxes has a total weight of 125 kilograms. How much does each type of box weigh?

OpenStudy (anonymous):

@GretaKnows

OpenStudy (anonymous):

@Studydiva10123

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (owlcoffee):

You can solve this by creating a system of equations. Let's call the "large boxes" a more suited name, "x" for example. And the small boxes "y", and we know the weight of each delivery and the amount of boxes used on such deliveries, so we will create a system of equations using that: \[1)2x + 3y = 56\] \[2) 4x + 8y = 125\] Now, it's just a matter of solving the system of equations.

OpenStudy (anonymous):

@Owlcoffee I will combine these two right?

OpenStudy (owlcoffee):

if you so desire, you can use whatever technique you deem fit.

OpenStudy (anonymous):

First of all let the large boxes to be (L) and small ones to be (S) Then from the part "A delivery of 2 large boxes and 3 small boxes has a total weight of 56 kilograms" U could conclude that : 2L + 3S = 56 lets denote this (eq.1) And from the part "A delivery of 4 large boxes and 8 small boxes has a total weight of 125 kilograms" U conclude that : 4L + 8S = 125 let's denote this (eq.2) So our objective nw is to solve this set of linear equations simultaneously. 2L + 3S = 56 & 4L + 8S = 125 So multiply (eq.1) by 2 and subtract it from (eq.2) you will get : 2S = 13 >> Then S = 13/2 which is known to be : 6.5 kg and then subs, by this value (S = 6.5) in any of two equations to get L : suppose we subs. in (eq.1): L = (56-3S)/2 >> L = 18.25 kg Let me know if you got it :)